# Effective-medium approach to the resonance distribution of wave scattering in a random point field

### v2

@@ III.2 Disorder-averaged Green function
Regarding the analytic properties of the average Green function at complex values of $k$, one notices the similarity between Eq. (39) and Eq. (9) for the free Green function $G(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$. Accordingly, Eq. (39) has two solutions valid for $k\in\mathbb{C}$: $\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ which exponentially vanishes for $\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|\rightarrow\infty$ on the domain $\operatorname{Im}k>0$, and $\langle Q^{-}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ which exponentially increases on the same domain.
Several additional comments should be made about the average Green function $\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$. First of all, if the medium is uniform and extends to infinity, then the average Green function is essentially the same as the free-space Green function
@@
where $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$, but with $k$ replaced by the effective wavenumber [28, 29, 63, 26, 25]
$\kappa(k)=\sqrt{k^{2}-nF(k)}\>.$ (41)
Due to the complex nature of the point scattering amplitude $F(k)$, the effective wavenumber $\kappa$ is itself complex, meaning that the average Green function exponentially decays with the distance $r$ from the point source.
Figure 2: Real part of the Green function in a 2D disk-shaped random point field for $N=300$ and $k=6\,\varsigma^{-1}$ with a source at the center ($\boldsymbol{\mathrm{r}}^{\prime}=\boldsymbol{\mathrm{0}}$). The solid curve is the numerical average over $2^{14}$ configurations, the dashed curve is the theoretical average (40), and the dotted curve is the free-space Green function (10) for comparison. The filled region is the numerically computed interquartile range of $\operatorname{Re}Q^{+}(k,r)$. The radius of the medium is indicated by the gray vertical line. The point scattering model (15) is used and the mean free path is $\ell_{\rm s}=1.5\,\varsigma$. Note that $\operatorname{Im}k\xrightarrow{>}0$ for all curves, but $\operatorname{Im}\kappa(k)>0$.
This decay is numerically confirmed with the program MSModel [56] in Fig. 2 for a two-dimensional (2D) point field. The agreement between the numerical average of $\operatorname{Re}Q^{+}(k,r)$ (solid) and the theoretical average $G^{+}(\kappa(k),r)$ (dashed) is so precise that the curves are completely superimposed. The exponential decay of $\langle Q^{+}(k,r)\rangle$ at large $r$ is evidenced by the comparison with the free-space Green function $G^{+}(k,r)$ (dotted). This supports the validity of the effective-medium equation (39).
Another important property is that the characteristic length of the exponential decay of the average Green function can be physically interpreted. If we take the square modulus of Eq. (40) using the asymptotic behavior (11), we get
@@
The imaginary part of $\kappa$, given by Eq. (41), can be expanded if we assume that $k^{2}$ is relatively large compared to $nF(k)$. We find
@@
Due to the optical theorem, given for instance by Eq. (44) of Ref. [46], it turns out that the last term of Eq. (43) is just the inverse of the scattering mean free path [25, 26, 64]
@@
The characteristic length $\ell_{\rm s}$ represents the mean distance between two successive collisions of the wave in the medium. This length plays a crucial role in transport theory. Moreover, since we consider complex wavenumbers in this work, we have ${k}_{\rm i}\neq 0$ in general. In anticipation of future calculations, we already introduce the notation
$\gamma=2{k}_{\rm i}\>.$ (45)
In this way, the behavior (42) of the average Green function reads
@@
The exponential decay of the form $\mathop{}\!\mathrm{e}^{-r/\ell_{\rm s}}$ in Eq. (46) is strongly reminiscent of the Beer-Lambert law for the attenuation (or extinction) of light in materials [65, 27]. Note, however, that the present decay actually corresponds to the loss of phase coherence of the wave [26, 25, 64, 52, *Doicu2019a2, *Doicu2019a3], because, due to the elastic collisions with the scatterers, the wave looses memory of its phase and thus of its direction of motion. As mentioned earlier, this decay does not indicate the presence of absorption of the wave by the medium.
It is worth noting that the mean free path $\ell_{\rm s}=(n\sigma)^{-1}$ is longer than the mean interscatterer distance $\varsigma=n^{-1/d}$. This results from the assumption that the point field is dilute, namely the size of the scatterers, which is determined by the scattering length $\alpha$ (defined as in Refs. [46, 47]), should be smaller than the interscatterer distance ($\alpha\ll\varsigma$). An important consequence of $\ell_{\rm s}\gg\varsigma$ is that the wave is likely to pass through many scatterers before a collision effectively occurs. This is not trivial because one could a priori think of the propagation of the wave as being simultaneously affected by all the scatterers. This supposition would lead to a characteristic distance of $\varsigma$ between successive interactions, or in other words $\ell_{\rm s}\simeq\varsigma$, which turns out to be incorrect, as pointed out in Ref. [25, p. 86].
Moreover, the presence of exponential evanescence in the Green function has an important consequence concerning the random and unpredictable character of the tracks formed by a quantum particle moving in a gaseous detector [44, 45]. Indeed, Eq. (46) strongly recalls the exponential probability density
$p_{\rm col}(x)\mathop{}\!\mathrm{d}x=\frac{1}{\ell_{\rm s}}\mathop{}\!\mathrm{e}^{-x/\ell_{\rm s}}\mathop{}\!\mathrm{d}x\>,$ (47)
that the next collision takes place at the distance $x$ from the previous collision point. Therefore, each time the wave encounters a scatterer, it gets a certain probability
$P_{\rm col}=1-\mathop{}\!\mathrm{e}^{-\varsigma/\ell_{\rm s}}\simeq\frac{\varsigma}{\ell_{\rm s}}\qquad\text{for}~{}\varsigma\ll\ell_{\rm s}\>,$ (48)
to collide with it. The probability (48) never reaches one, except of course in the nonphysical limit $\ell_{\rm s}=0$. This shows that, even in quantum mechanics, the collisions constitute a Poisson process. This is an important observation because it supports the idea that the visible track of a quantum particle, such as an electron, in a gaseous detector is random and unpredictable.
III.3 Low-energy peaks in the resonance density
An important application of the effective-medium equation (39) is to provide an explanation for the low-energy peaks in the resonance density discovered in our previous paper [47]. As discussed in Sec. II.3, the resonances appear as poles of the full Green function $Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ in the lower half-plane of the wavenumber ($\operatorname{Im}k<0$). In order to facilitate the comparison between theory and simulation, we restrict ourselves to the case of a spherical-shaped disordered medium of radius $R$, and we solve the effective-medium equation (39) for the resonance poles. This calculation is shown in details in Appendix B using the partial wave method [57, 58, 59, 60]. The result is the equation
for a complex resonance at $k\in\mathbb{C}$. In Eq. (49), $\kappa$ is the effective wavenumber (41), $J_{\nu}(z)$ and $H^{+}_{\nu}(z)$ are the Bessel and Hankel functions, respectively, and the order $\nu$ is given by
$\nu=\ell+\frac{d-2}{2}\>,$ (50)
where $\ell$ is the orbital quantum number.
The comparison between the numerical solutions of Eq. (49) for $k\in\mathbb{C}$ and the actual peaks of the resonance density obtained with the program MSModel [56] is shown in Fig. 3(a) for a 2D point field and in Fig. 3(b) for a three-dimensional (3D) point field.
Figure 3: (a) Resonance density $2\pi\varrho^{(2)}({k}_{\rm r},{k}_{\rm i})$ computed with the program MSModel [56] in a 2D disk-shaped random point field for $N=50$ and using the point scattering model given by Eq. (18) of Ref. [47] with $\alpha=0.1\,\varsigma$. The density is averaged over $2^{8}$ random configurations. The resonances of the effective medium, obtained by numerically solving Eq. (49), are depicted by small circles labelled with the corresponding angular momentum $\ell$. The dashed line separates the low from the high energy region (52). (b) Same as panel (a) but in a 3D spherical random point field for $N=100$ averaged over $2^{8}$ configurations.
In these two panels, instead of Eq. (15), we exceptionally use the point scattering model given by Eq. (18) of Ref. [47] with the scattering length $\alpha=0.1\,\varsigma$. Indeed, this point scattering model is more appropriate than Eq. (15) to explore the small-$k$ regime because the cross section does not strongly diverge for $k\rightarrow 0$.
The adequacy between the roots of Eq. (49) and the peaks in Figs. 3(a) and 3(b) is remarkable and supports the validity of the effective-medium equation (38) at large wavelength ($k\varsigma\ll 1$). It shows that, in this low-energy regime, the incident wave mainly perceives the random point field as a semitransparent sphere without the details due to disorder.
Moreover, it turns out that the resonance structure shown in Fig. 3 is little affected by the geometric shape of the disordered medium or the scattering parameters such as the cross section of the point scatterers. Indeed, the same kind of structure can be observed for a cubic medium, instead of a spherical medium. The robustness of this resonance structure allows us to make a quite radical simplification of the resonance equation (49). Indeed, given the weak influence of the cross section of the point scatterers, it is possible to let the parameter $\kappa$ tend to infinity, leading to
$H^{+}_{\nu}(kR)=0\>.$ (51)
As explained in Appendix C, this is the equation of resonances corresponding to the scattering by a hard sphere of radius $R$. The roots of Eq. (51) are shown in the figure of Appendix C and indeed form structures very close to those in Fig. 3. This observation confirms the robustness of this resonance structure against possibly significant changes of the scattering parameters.
Nevertheless, the hard-sphere resonance equation (51) is not as accurate as Eq. (49) because it neglects the scattering within the medium. This is why the roots of Eq. (51) are not shown in Fig. 3. In contrast, the prediction of the resonance equation (49) of the effective-medium theory is more reliable.
In the 2D case [Fig. 3(a)], a slight shift between simulation and prediction can be noticed. In addition, the predictions in the upper right side of Fig. 3(a) do not match the resonance band close to the real $k$ axis. These disagreements could be due to an effect of Anderson localization because, in 2D, the point cross section is not bounded at low energy due to a logarithmic singularity at $k=0$. Therefore, the product $k\ell_{\rm s}=\tfrac{k}{n\sigma(k)}$ becomes small compared to 1, which thus favors Anderson localization in a finite 2D system [31, 66, 26, 32, 64, 15].
On another side, the agreement between the theoretical estimate (49) and the numerical simulation is much more striking in the 3D case [Fig. 3(b)]. Even better: the theoretical prediction coincides with the faint peaks closest to the real axis for $\ell=0,1,2$ in Fig. 3(b). This is not obvious because these peaks constitute the leftmost edge of the resonance band which was not expected to be described by this theory. At higher energies, however, that is when
$\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}\>,$ (52)
the prediction (49) does no longer match the position of the resonance band in the upper right side of Fig. 3(b). In order to describe this band, we need to consider a more complete transport theory which remains valid at wavelengths smaller than the mean interscatterer distance ($\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$). This is the subject of Sec. IV.
IV Semiclassical transport
The issue with the effective-medium approach is that, due to the destructive interference caused by the disorder averaging, only the first collision is taken into account. When the quantum particle propagates further in the medium, multiple scattering becomes more significant. This explain why the effective-medium equation (39) is unable to describe the resonance band in Fig. 3 in the high-energy region ($\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$). To address this issue, we need to consider a more elaborate theory for wave transport.
In this section, we introduce wave transport theory and compare its prediction of the disorder-averaged square modulus of the wavefunction, $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$, to the results of numerical simulations based on the Foldy-Lax model. In particular, we want to know how good the prediction of $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$ is compared to a typical realization of $\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}$ for a single configuration of the disorder in the regime of large system ($N\gg 1$). To this end, we let the quantum particle start from the center of the spherical medium.
Finally, we show that the wave transport theory gives an accurate prediction of the maximum of the resonance band in the high-energy region.
Throughout this section, we assume $\hbar=2m=1$ for convenience, so that the energy is also denoted as $E=\hbar\omega=\omega$, the energy-momentum relation becomes $\omega=k^{2}$, and the group velocity reads $v=2k$.
IV.1 Time evolution and complex frequencies
Besides the wavefunction, a relevant quantity in quantum mechanics is the density matrix defined for a pure state $\left|\psi(t)\right\rangle$ by
$\hat{\rho}(t)=\left|\psi(t)\right\rangle\left\langle\psi(t)\right|\>.$ (53)
This matrix contains much more information about the particle state than the probability distribution alone, especially the momentum, the direction of motion and the coherence, given by its off-diagonal elements. Moreover, the probability density is just the diagonal part of the density matrix: $\rho(\boldsymbol{\mathrm{r}},t)=\left\langle\boldsymbol{\mathrm{r}}\right|\hat{\rho}(t)\left|\boldsymbol{\mathrm{r}}\right\rangle$. The time evolution of the density matrix is governed by the quantum Liouville equation
$\mathrm{i}\frac{\partial\hat{\rho}(t)}{\partial t}=[\hat{H},\hat{\rho}(t)]\>,$ (54)
where $\hat{H}=\hat{\boldsymbol{\mathrm{p}}}^{2}+U(\hat{\boldsymbol{\mathrm{r}}})$ is the Hamiltonian of the multiple scattering problem. The initial condition of Eq. (54) at $t=0$ is $\hat{\rho}_{0}$. In the following sections, we will typically assume that the initial state is the position eigenstate
@@ IV.1 Time evolution and complex frequencies
$\hat{\rho}_{0}=\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|\>,$ (55)
but the present calculation is much more general. Since the Hamiltonian is time-independent, the solution of Eq. (54) simply reads
$\hat{\rho}(t)=\mathop{}\!\mathrm{e}^{-\mathrm{i}\hat{H}t}\hat{\rho}_{0}\mathop{}\!\mathrm{e}^{\mathrm{i}\hat{H}t}\>.$ (56)
The point is that, instead of calculating $\hat{\rho}(t)$ as a time-dependent function, it turns out to be more convenient to determine the Laplace transform of $\hat{\rho}(t)$ in time
$\hat{\rho}(s)=\int_{0}^{\infty}\mathop{}\!\mathrm{e}^{-st}\hat{\rho}(t)\mathop{}\!\mathrm{d}t\>.$ (57)
Another way is to use the Fourier transform of $\hat{\rho}(t)$, as usually done in the literature [25, 26]. However, we find the Laplace transform more suitable than the Fourier transform for non-oscillating signals such as the density $\rho(\boldsymbol{\mathrm{r}},t)$. The advantage of the Laplace transform $\hat{\rho}(s)$ is that it contains the same information as the time-dependent density matrix $\hat{\rho}(t)$, while being a stationary quantity.
In order to evaluate (57), we represent the evolution operators in the frequency space
$\mathop{}\!\mathrm{e}^{\pm\mathrm{i}\hat{H}t}=\int_{\mathbb{R}\mp\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}\mathop{}\!\mathrm{e}^{\pm\mathrm{i}\omega t}\hat{Q}(\omega)\>,$ (58)
where $\hat{Q}(\omega)$ is the resolvent of the full problem
$\hat{Q}(\omega)=\frac{1}{\omega-\hat{H}}\qquad\forall\omega\in\mathbb{C}\setminus\mathbb{R}^{+}\>,$ (59)
introduced in Eq. (26). The parameter $\varepsilon\xrightarrow{>}0$ in Eq. (58) is an infinitesimal shift of the integration contour in the imaginary direction. This is required for convergence due to the singularities in the resolvent (59). Substituting Eq. (56) and (58) into Eq. (57) yields
$\begin{split}\hat{\rho}(s)&=\int_{\mathbb{R}+\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}\int_{\mathbb{R}-\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega^{\prime}}{2\pi}\\ &\times\int_{0}^{\infty}\mathop{}\!\mathrm{d}t\mathop{}\!\mathrm{e}^{-(s+\mathrm{i}\omega-\mathrm{i}\omega^{\prime})t}\hat{Q}(\omega)\hat{\rho}_{0}\hat{Q}(\omega^{\prime})\>.\end{split}$ (60)
Integrating over the time $t$, and using the changes of variable
$\omega_{\rm m}=\frac{\omega+\omega^{\prime}}{2}\>,\qquad\omega_{\rm d}=\omega-\omega^{\prime}\>,$ (61)
Eq. (60) becomes
$\hat{\rho}(s)=\int_{\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\omega_{\rm m}}{2\pi}\int_{\mathbb{R}+\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega_{\rm d}}{2\pi}\>\frac{\hat{Q}(\omega_{\rm m}+\tfrac{\omega_{\rm d}}{2})\hat{\rho}_{0}\hat{Q}(\omega_{\rm m}-\tfrac{\omega_{\rm d}}{2})}{s+\mathrm{i}\omega_{\rm d}}\>.$ (62)
Finally, the frequency difference, $\omega_{\rm d}$, can be integrated out with the residue theorem around the simple pole at $\omega_{\rm d}=\mathrm{i}s$. The result is
$\hat{\rho}(s)=\int_{\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\omega_{\rm m}}{2\pi}\>\hat{Q}(\omega_{\rm m}+\tfrac{\mathrm{i}}{2}s)\hat{\rho}_{0}\hat{Q}(\omega_{\rm m}-\tfrac{\mathrm{i}}{2}s)\>.$ (63)
To get rid of the remaining integral, we assume that the spectrum of the state $\hat{\rho}_{0}$ is very peaked around the central frequency $\omega_{0}$. Of course, one could argue that the monochromatic assumption of the state $\hat{\rho}_{0}$ is not compatible with the initial condition that $\hat{\rho}_{0}=\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|$, because obviously the mean frequency of $\hat{\rho}_{0}$ cannot equal $\omega_{0}$, i.e., $\operatorname{Tr}(\hat{\rho}_{0}\hat{H})\neq\omega_{0}$, due to the uncertainty principle. However, it is possible to overcome this issue by treating $\boldsymbol{\mathrm{r}}_{0}$ as a source point for a wave at a central frequency $\omega_{0}$ during a time window of duration $\Delta t$ such that $s\ll\tfrac{2\pi}{\Delta t}\ll\omega_{0}$, as presented for instance in Ref. [67, Sec. 4.2.1]. The result is the reduction of the density matrix (63) to
$\hat{\rho}(s)=\hat{Q}(\omega_{0}+\tfrac{\mathrm{i}}{2}s)\hat{\rho}_{0}\hat{Q}(\omega_{0}-\tfrac{\mathrm{i}}{2}s)\>.$ (64)
This is the sought expression for the Laplace transform of the density matrix. It is showed in Appendix D that the normalization of the density matrix, $\operatorname{Tr}\hat{\rho}(t)$, is conserved over time, as it should in absence of absorption. This appendix also provides the absolute normalization of $\hat{\rho}(t)$ in Eq. (64), which is not 11 11.
We notice on Eq. (64) the other advantage of the Laplace transform on the usual Fourier transform [25, 26]. Indeed, the right-hand Green function in Eq. (64) is just the Hermitian conjugate of the left-hand Green function, according to $\hat{Q}^{\dagger}(\omega)=\hat{Q}({\omega}^{*})$. Therefore, the Green function only has to be evaluated at the single complex frequency $\omega_{+}=\omega_{0}+\tfrac{\mathrm{i}}{2}s$. If we had calculated the Fourier transform, $s$ would have been imaginary, leading to two different real frequencies in Eq. (64), and thus less convenient expressions. This shows the interest of the Laplace transform in this context.
The result (64) shows that the Laplace variable $s$ can be interpreted as the imaginary part of the frequency $\omega$. It is an important result because it proves the equivalence between the evolution of the density matrix in the time domain and the complex frequency domain. In addition, Eq. (64) plays a crucial role in this paper because it relates $\hat{\rho}(s)$ to the spectrum of complex resonances given by the Green operator $\hat{Q}(\omega)$.
The density matrix (64) is likely more transparent when expressed in position space. Using the initial condition (55), the density matrix reads [25, 26]
$\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime},s)=Q^{+}(k_{+},\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})Q^{-}(k_{-},\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{r}}_{0})\>,$ (65)
or equivalently,
$\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime},s)=Q^{+}(k_{+},\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0}){[Q^{+}(k_{+},\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{r}}_{0})]}^{*}\>.$ (66)
The wavenumbers $k_{\pm}$ in Eqs. (65)–(66) corresponding to the frequencies $\omega_{\pm}=\omega_{0}\pm\tfrac{\mathrm{i}}{2}s$ are given by
$k_{\pm}=k_{0}\pm\mathrm{i}\frac{s}{2v}\>,$ (67)
for the group velocity $v$. This also shows that $s=\gamma v$ according to the notation (45). Appendix E presents a consistency check that, in free space, Eq. (66) can indeed be interpreted as a freely propagating wave packet in the time domain.
IV.2 Semiclassical transport equation
It is possible to obtain a closed equation for the density matrix (66) after averaging over the random configurations of the scatterers, as shown in Appendix F. This leads to an integral transport equation of the same form as the Bethe-Salpeter equation [28, 68, 19, 20, 21, 24, 25, 26, 69]
$\begin{split}\left\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime},s)\right\rangle&=\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\langle{Q^{+}(k,\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{r}}_{0})}^{*}\rangle\\ &+\int_{\mathbb{R}^{d}}\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}})\rangle\langle{Q^{+}(k,\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{x}})}^{*}\rangle\\ &\qquad\times\left|F(k)\right|^{2}\left\langle\rho(\boldsymbol{\mathrm{x}},\boldsymbol{\mathrm{x}},s)\right\rangle n(\boldsymbol{\mathrm{x}})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}\>,\end{split}$ (68)
where the wavenumber has the complex value $k=k_{+}$ according to Eq. (67). Equation (68) describes the multiple scattering of the particle in the disordered medium. The last term of the right-hand side accounts for the individual collisions which may occur anywhere in the medium, hence the integral over $\boldsymbol{\mathrm{x}}$.
@@ IV.2 Semiclassical transport equation
The main assumption behind the Bethe-Salpeter equation (68) is that the wavelength is much smaller than the mean free path [^2] Note that, since the mean free path is at least equal to the mean interatomic distance ($\ell_{\rm s}=\varsigma$), the condition (69) is necessarily satisfied for $k_{0}\varsigma\gg 1$.
@@
This is indeed equivalent to $\omega_{0}\gg n\sigma v$ as long as the dispersion relation is a power law: $\omega(k)\propto k^{a}~{}(\forall a>0)$. The condition (69) is known in the literature as the weak scattering regime [26], or the weak disorder regime [25]. In this regime, coherent effects such as localization play less significant roles, especially in 2D and 3D. Note that, in the context of gaseous particle detectors, the condition (69) is usually met. In this regard, some orders of magnitude are shown in Tab. 1.

### v3

@@ II.3 Disorder-averaged Green function
where $n(\boldsymbol{\mathrm{r}})$ is the scatterer density (5) and $F(k)$ is the point scattering amplitude (12). Equation (30) is very close to the Green-function equation associated to Eq. (2) but with the potential $U(\boldsymbol{\mathrm{r}})$ replaced by $n(\boldsymbol{\mathrm{r}})F(k)$. By analogy with Eq. (9), the average Green function has two solutions valid for $k\in\mathbb{C}$: $\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ which exponentially vanishes for $\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|\rightarrow\infty$ on the domain $\operatorname{Im}k>0$, and $\langle Q^{-}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ which exponentially increases on the same domain.
Often, it is convenient to introduce the (local) effective wavenumber [28, 29, 64, 26, 27]
$\kappa(k,\boldsymbol{\mathrm{r}})=\sqrt{k^{2}-n(\boldsymbol{\mathrm{r}})F(k)}\>,$ (31)
such that, in regions of constant $n(\boldsymbol{\mathrm{r}})$, the average Green function essentially behaves as the free Green function
@@
where $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$. Due to the complex nature of the point scattering amplitude $F(k)$, the effective wavenumber $\kappa$ is itself complex, meaning that the average Green function exponentially decays with the distance $r$ from the point source, even for $k\in\mathbb{R}$ because of disorder. This decay can be highlighted using the asymptotic behavior (11):
@@
The imaginary part of $\kappa$, given by Eq. (31) with $n(\boldsymbol{\mathrm{r}})=n$, can be expanded if we assume that $k^{2}$ is relatively large compared to $nF(k)$. We find
@@
In virtue of the optical theorem, given for instance by Eq. (44) of Ref. [46], it turns out that the last term of Eq. (34) is just the inverse of the scattering mean free path [26, 27, 65]
@@
The characteristic length $\ell_{\rm s}$ represents the mean distance between two successive collisions of the wave in the medium and thus plays a crucial role in transport theory. This decay represents the loss of coherence of the incident wave due to the disorder. Another contribution to the decay in Eq. (34) comes from the intrinsic imaginary part of $k$ which is essential for the study of complex resonances as we saw in Eq. (25). In anticipation of future calculations, we already introduce the notation
$\gamma=2{k}_{\rm i}\>,$ (36)
such that the behavior (33) of the average Green function reads
@@
We will see in Sec. III.1 that the effective-medium equation (30) can be used to predict the peaks in the resonance density in the low-energy region.
II.4 Semiclassical transport
The issue with the effective-medium equation (30) is that, due to the destructive interference caused by the disorder averaging, the wave undergoes exponential damping at the scale of the mean free path, alike in absorption processes. Therefore, this equation is not adapted to describe the multiple scattering which takes place on scales larger than one mean free path and which is responsible in particular for the spatial diffusion of the wave intensity. We may expect that this multiple scattering will significantly affect the distribution of the complex resonances since their location is governed by the characteristic escape rate from the disordered region. To address the issue of multiple scattering, we need to consider a more elaborate theory for wave transport that is referred to as the semiclassical theory [66, 67]. The central quantity of this theory is the intensity of the full Green function
$\rho(\boldsymbol{\mathrm{r}},\gamma)=\left|Q^{+}(k_{0}+\tfrac{\mathrm{i}}{2}\gamma,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\right|^{2}\>,$ (38)
where $\boldsymbol{\mathrm{r}}_{0}$ is the source point of the wave. Note that $\rho(\boldsymbol{\mathrm{r}},\gamma)$ fluctuates with the disorder as seen in Fig. 1(a) but possesses a well-defined average shown in Fig. 1(b). The equation governing the disorder-averaged intensity, $\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\rangle$, is known as the Bethe-Salpeter equation and reads [28, 68, 19, 20, 21, 24, 26, 27, 69, 62, 63]
@@ II.4 Semiclassical transport
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})+\int_{\mathbb{R}^{d}}\mathop{}\!\mathrm{d}{\boldsymbol{\mathrm{r}}^{\prime}}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})n(\boldsymbol{\mathrm{r}}^{\prime})\sigma(k)\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\>,$ (39)
where the transport kernel $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ is defined by the square modulus of the effective Green function
$K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\xrightarrow{r\rightarrow\infty}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}\>.$ (41)
Note that the small imaginary part of $k=k_{0}+\mathrm{i}{k}_{\rm i}$ can be neglected in the cross section because the equation only considers time scales much longer than the oscillation period of the wave (${k}_{\rm i}\ll k_{0}$). It is thus safe to write $\sigma(k)\simeq\sigma(k_{0})$.
@@
The main assumption behind the Bethe-Salpeter equation (39) is that the wavelength is much smaller than the mean free path [^1] Note that, since the mean free path is at least equal to the mean interatomic distance ($\ell_{\rm s}\geq\varsigma$), the condition (42) is necessarily satisfied for $k_{0}\varsigma\gg 1$.
@@
The condition (42) is known in the literature as the weak scattering regime [26], or the weak disorder regime [27]. In this regime, coherent effects such as localization play less significant roles, especially in 2D and 3D. Note that, in the context of gaseous particle detectors, the condition (42) is usually met. In this regard, some orders of magnitude are given in Tab. 1.